# Efficient Evaluation of LLM Performance with Statistical Guarantees

### v1

@@ Abstract
Exhaustively evaluating many large language models (LLMs) on a large suite of benchmarks is expensive. We cast benchmarking as finite-population inference and, under a fixed query budget, seek tight confidence intervals (CIs) for model accuracy with valid frequentist coverage. We propose Factorized Active Querying (FAQ), which (a) leverages historical information through a Bayesian factor model; (b) adaptively selects questions using a hybrid variance-reduction/active-learning sampling policy; and (c) maintains validity through Proactive Active Inference—a finite-population extension of active inference (Zrnic & Candes, 2024) that enables direct question selection while preserving coverage. With negligible overhead cost, FAQ delivers up to $5\times$ effective sample size gains over strong baselines on two benchmark suites, across varying historical-data missingness levels: this means that it matches the CI width of uniform sampling while using up to $5\times$ fewer queries. We release our source code and our curated datasets to support reproducible evaluation and future research.
@@ 1 Introduction
A common practice is to evaluate candidate models’ performances on a finite, curated benchmark suite. Such suites increasingly include questions that (a) require expensive domain-expert adjudication (often by multiple annotators) (Szymanski et al., 2025; Li et al., 2025; Enguehard et al., 2025; Croxford et al., 2025); (b) involve multi-turn interactions (e.g., helpfulness and safety) (He et al., 2024; Kwan et al., 2024); and/or (c) demand long-form reasoning with verification (Bai et al., 2024; Ding et al., 2025). With tens of thousands of questions (e.g., 1212K on MMLU-Pro (Wang et al., 2024)), and hundreds to thousands of candidate models, exhaustively evaluating every model on every question is expensive in API costs, runtime, and expert annotation (Zhou et al., 2025). In practice, organizations may instead resort to “vibes”-based judgments on ad-hoc subsets (Mazzoni & Sharma, 2025).
@@
    A coverage-preserving finite-population extension of active inference (Zrnic & Candes, 2024), which we call Proactive Active Inference (PAI). Unlike active inference, PAI directly selects the next question to query in a data-adaptive manner, rather than rigidly proceeding according to a fixed ordering and only making label/skip decisions. [^1] While we focus on finite-bank benchmarking, PAI is a general coverage-preserving framework for adaptive querying over finite populations: we will explore its broader potential in future work.
@@ 2 Related Work
Probabilistic factor modeling    FAQ is inspired by probabilistic factor modeling (Mnih & Salakhutdinov, 2007; Collins et al., 2001; Chen et al., 2023), which assumes a (Bayesian) generative factor model for a partially-observed binary matrix—here, the matrix of model-question correctness outcomes—and learns low-dimensional latent factors from the data. Importantly, any uncertainty guarantees for these approaches are tied to correct specification of the factor model (and any prior assumptions). In contrast, FAQ uses the Bayesian factor model only as a mental model to guide adaptive query selection, while delivering valid frequentist coverage regardless of factor-model fidelity. If the data is well-described by the factor model, FAQ automatically adapts and delivers tighter CIs.
Active statistical inference    For model-free frequentist coverage, FAQ builds on active statistical inference (Zrnic & Candes, 2024; Gligorić et al., 2025), which uses machine learning predictions over unlabeled covariates/features to adaptively select labels under a fixed budget and construct CIs. Our benchmark setting, however, provides no explicit covariates, so we use the factor model above to construct latent features to guide adaptive query selection. Moreover, unlike sequential active inference—which, in a pre-specified “nature’s order” stream, can only make label/skip decisions for the next item and enforces the budget constraints only in expectation—PAI/FAQ directly selects the next question from the full benchmark bank using all available information and enforces the query budget exactly.
@@
Active learning    FAQ draws on Bayesian active learning ideas (c.f. e.g., Settles, 2009; Di Fiore et al., 2023; Cohn et al., 1994; MacKay, 1992; Fedorov, 2013; Schein & Ungar, 2007) to guide question selection. One component of our hybrid policy upweights questions predicted (under the factor model) to yield larger posterior variance reductions for $\theta$. Because our goal is tight frequentist uncertainty quantification of $\theta$, we combine this active-learning component with an oracle-inspired variance-reduction component.
@@ 3.2 Proactive Active Inference
Sequential active inference (Zrnic & Candes, 2024) provides coverage, but differs from our approach in two key ways. First, it follows nature’s order: at time $t$, one observes only the next item in a fixed stream and can only make a label/skip decision via an adaptive Bernoulli policy. In contrast, in our benchmark-bank setting, we have access to the full pool of question factors $\{v_{j}\}_{j=1}^{N_{q}}$ at every step and wish to leverage this global information to proactively select the next question. Second, we desire a workflow that is deterministic conditional on selection: once a question is chosen, it is labeled with probability $1$ (no coin flip).
@@ 4 Experimental Setup
with baseline predictor $p_{\text{base}}(t)\in\{0,\overline{p}_{t}\}$. We consider: (a) $\pi_{t}(t)=n_{b}/N_{q}$, (Bernoulli); (b) $\pi_{t}(t)\propto\sqrt{\overline{p}_{t}(1-\overline{p}_{t})}$ (Neyman); and (c) $\pi_{t}(t)\propto\min(\overline{p}_{t},1-\overline{p}_{t})$ (c.f. Zrnic & Candes (2024)). For (b)-(c), we normalize to probabilities, apply $\tau$-mixing, and use the recommended budget-stabilization wrapper in Equation 8 of Zrnic & Candes (2024). Note that the choice of $p_{\text{base}}(t)=0$ and $\pi_{t}(t)=n_{b}/N_{q}$ reduces to uniform sampling.
Traditional active inference ablation    Since the baselines above omit the factor model—a core component of FAQ—we additionally ablate FAQ to isolate the contribution of PAI and our factor-model-guided hybrid sampling policy. Specifically, we replace PAI and hybrid sampling with Zrnic & Candes (2024)’s sequential active inference—keeping the same factor-model predictions $\hat{p}_{j}^{(t)}$, using their recommended labeling policies (see Eqs. 3 and 8 therein), and tuning their $\tau$-mixing parameter (Appendix C).
@@ 5 Empirical Results
Methodological gains    To isolate the gains of PAI and our hybrid sampling (beyond the factor model), we compare FAQ to a “traditional” active inference ablation that follows Zrnic & Candes (2024)’s sequential (“nature’s order”) [^12] Here, “nature’s order” refers to the benchmarks’ native question orderings in HuggingFace’s Open LLM Leaderboard. procedure while using the same factor-model predictions. We report CI-width ratios (rather than ESS) since they compare tightness directly across budgets without converting to a uniform-equivalent “budget savings” scale. Figure 5 reveals a clear phase shift. At low-budgets (e.g, $10\%$ of the bank), FAQ yields substantially narrower CIs; at larger budgets, the traditional ablation eventually overtakes FAQ by similar margins. This is not discouraging: large-scale evaluation (11K+ models, 1010K+ questions) typically emphasizes the low-budget regime where FAQ is strongest. It also suggests a natural extension—a three-stage hybrid policy: active-learning early, oracle-style variance reduction mid-budget, and traditional active inference at the end—combining the best of both worlds.
@@ A.2 Derivation of Online Logistic Regression Laplace Approximation Update
This derivation was inspired by Ch. 4 of (Bishop & Nasrabadi, 2006)’s discussion on Laplace approximations and Bayesian logistic regression. The key idea is to Taylor expand the log-posterior around the current mean, which yields approximate closed-form update formulas while preserving approximate Gaussian conjugacy.
@@ A.3.2 Technical Assumptions/Conditions
Assumptions A.1-A.2 are the standard conditions under which a CLT for martingale difference arrays holds (c.f. Hall & Heyde (2014, Corollary 3.2)). Assumption A.3 is mild and is used to facilitate consistency of the variance estimator. Intuitively, in our setting of LLM evaluation with our proposed hybrid sampling policy (c.f. Eq. 10), the variance stabilization assumption is reasonable because as more outcome labels are observed, the model-factor estimate $(\hat{u}^{(t)},\hat{\Sigma}^{(t)})$ stabilizes, which stabilizes $\hat{p}^{(t)}$, and hence $q_{t}(\cdot)$, too (because $(\alpha_{t},\beta_{t})$ also stabilize). In addition, the Lindeberg condition is satisfied, by construction, under our hybrid sampling policy because $\tau$-mixing lower-bounds $q_{t}(j)$, avoiding extreme inverse-probability weights, and since $z_{j},\hat{p}^{(t)}_{j}\in[0,1]$, the increments $\Delta_{t}$ are uniformly bounded (by a constant depending only on $\tau$). The $\tau$-mixing also ensures the satisfaction of Assumption A.3.
@@ A.3.3 Asymptotic Normality via Martingale CLT
Under Assumptions A.1-A.2, the martingale central limit theorem (c.f.,  Hall & Heyde (2014, Corollary 3.2)) implies
@@
    Define $\tilde{\theta}^{(t-1)}:=\frac{1}{N_{q}}\sum_{j=1}^{N_{q}}\hat{p}^{(t-1)}_{j}$. Under Assumptions A.1A.2, Hall & Heyde (2014, Theorem 2.23) guarantees that
@@
    To show (17), we apply the Lenglart inequality (c.f., e.g., Whitt (2007, Lemma 3.7) or Jacod & Protter (2012, pp. 45)) to obtain

### v2

@@ Abstract
Exhaustively evaluating many large language models (LLMs) on a large suite of benchmarks is expensive. We cast benchmarking as finite-population inference and, under a fixed query budget, seek tight confidence intervals (CIs) for model accuracy with valid frequentist coverage. We propose Factorized Active Querying (FAQ), which (a) leverages historical information through a Bayesian factor model; (b) adaptively selects questions using a hybrid variance-reduction/active-learning sampling policy; and (c) maintains validity through Proactive Active Inference—a finite-population extension of active inference (Zrnic and Candès, 2024) that enables direct question selection while preserving coverage. With negligible overhead cost, FAQ delivers up to $5\times$ effective sample size gains over strong baselines on two benchmark suites, across varying historical-data missingness levels: this means that it matches the CI width of uniform sampling while using up to $5\times$ fewer queries. We release our source code and our curated datasets to support reproducible evaluation and future research.
@@ 1 Introduction
A common practice is to evaluate candidate models’ performances on a finite, curated benchmark suite. Such suites increasingly include questions that (a) require expensive domain-expert adjudication (often by multiple annotators) (Szymanski et al., 2025; Li et al., 2025; Enguehard et al., 2025; Croxford et al., 2025); (b) involve multi-turn interactions (e.g., helpfulness and safety) (He et al., 2024; Kwan et al., 2024); and/or (c) demand long-form reasoning with verification (Bai et al., 2024; Ding et al., 2025). With tens of thousands of questions (e.g., 1212K on MMLU-Pro (Wang et al., 2024)), and hundreds to thousands of candidate models, exhaustively evaluating every model on every question is expensive in API costs, runtime, and expert annotation (Zhou et al., 2025). In practice, organizations may instead resort to “vibes”-based judgments on ad-hoc subsets (Mazzoni and Sharma, 2025).
@@
    A coverage-preserving finite-population extension of active inference (Zrnic and Candès, 2024), which we call Proactive Active Inference (PAI). Unlike active inference, PAI directly selects the next question to query in a data-adaptive manner, rather than rigidly proceeding according to a fixed ordering and only making label/skip decisions. [^1] While we focus on finite-bank benchmarking, PAI is a general coverage-preserving framework for adaptive querying over finite populations: we will explore its broader potential in future work.
@@ 2 Related Work
Probabilistic factor modeling    FAQ is inspired by probabilistic factor modeling (Mnih and Salakhutdinov, 2007; Collins et al., 2001; Chen et al., 2023), which assumes a (Bayesian) generative factor model for a partially-observed binary matrix—here, the matrix of model-question correctness outcomes—and learns low-dimensional latent factors from the data. Importantly, any uncertainty guarantees for these approaches are tied to correct specification of the factor model (and any prior assumptions). In contrast, FAQ uses the Bayesian factor model only as a mental model to guide adaptive query selection, while delivering valid frequentist coverage regardless of factor-model fidelity. If the data is well-described by the factor model, FAQ automatically adapts and delivers tighter CIs.
Active statistical inference    For model-free frequentist coverage, FAQ builds on active statistical inference (Zrnic and Candès, 2024; Gligorić et al., 2025), which uses machine learning predictions over unlabeled covariates/features to adaptively select labels under a fixed budget and construct CIs. Our benchmark setting, however, provides no explicit covariates, so we use the factor model above to construct latent features to guide adaptive query selection. Moreover, unlike sequential active inference—which, in a pre-specified “nature’s order” stream, can only make label/skip decisions for the next item and enforces the budget constraints only in expectation—PAI/FAQ directly selects the next question from the full benchmark bank using all available information and enforces the query budget exactly.
@@
Active learning    FAQ draws on Bayesian active learning ideas (c.f. e.g., Settles, 2009; Di Fiore et al., 2023; Cohn et al., 1994; MacKay, 1992; Fedorov, 2013; Schein and Ungar, 2007) to guide question selection. One component of our hybrid policy upweights questions predicted (under the factor model) to yield larger posterior variance reductions for $\theta$. Because our goal is tight frequentist uncertainty quantification of $\theta$, we combine this active-learning component with an oracle-inspired variance-reduction component.
@@ 3.2 Proactive Active Inference
Sequential active inference (Zrnic and Candès, 2024) provides coverage, but differs from our approach in two key ways. First, it follows nature’s order: at time $t$, one observes only the next item in a fixed stream and can only make a label/skip decision via an adaptive Bernoulli policy. In contrast, in our benchmark-bank setting, we have access to the full pool of question factors $\{v_{j}\}_{j=1}^{N_{q}}$ at every step and wish to leverage this global information to proactively select the next question. Second, we desire a workflow that is deterministic conditional on selection: once a question is chosen, it is labeled with probability $1$ (no coin flip).
@@ 4 Experimental Setup
with baseline predictor $p_{\text{base}}(t)\in\{0,\overline{p}_{t}\}$. We consider: (a) $\pi_{t}(t)=n_{b}/N_{q}$, (Bernoulli); (b) $\pi_{t}(t)\propto\sqrt{\overline{p}_{t}(1-\overline{p}_{t})}$ (Neyman); and (c) $\pi_{t}(t)\propto\min(\overline{p}_{t},1-\overline{p}_{t})$ (c.f. Zrnic and Candès (2024)). For (b)-(c), we normalize to probabilities, apply $\tau$-mixing, and use the recommended budget-stabilization wrapper in Equation 8 of Zrnic and Candès (2024). Note that the choice of $p_{\text{base}}(t)=0$ and $\pi_{t}(t)=n_{b}/N_{q}$ reduces to uniform sampling.
Traditional active inference ablation    Since the baselines above omit the factor model—a core component of FAQ—we additionally ablate FAQ to isolate the contribution of PAI and our factor-model-guided hybrid sampling policy. Specifically, we replace PAI and hybrid sampling with Zrnic and Candès (2024)’s sequential active inference—keeping the same factor-model predictions $\hat{p}_{j}^{(t)}$, using their recommended labeling policies (see Eqs. 3 and 8 therein), and tuning their $\tau$-mixing parameter (Appendix C).
@@ 5 Empirical Results
Methodological gains    To isolate the gains of PAI and our hybrid sampling (beyond the factor model), we compare FAQ to a “traditional” active inference ablation that follows Zrnic and Candès (2024)’s sequential (“nature’s order”) [^12] Here, “nature’s order” refers to the benchmarks’ native question orderings in HuggingFace’s Open LLM Leaderboard. procedure while using the same factor-model predictions. We report CI-width ratios (rather than ESS) since they compare tightness directly across budgets without converting to a uniform-equivalent “budget savings” scale. Figure 5 reveals a clear phase shift. At low-budgets (e.g, $10\%$ of the bank), FAQ yields substantially narrower CIs; at larger budgets, the traditional ablation eventually overtakes FAQ by similar margins. This is not discouraging: large-scale evaluation (11K+ models, 1010K+ questions) typically emphasizes the low-budget regime where FAQ is strongest. It also suggests a natural extension—a three-stage hybrid policy: active-learning early, oracle-style variance reduction mid-budget, and traditional active inference at the end—combining the best of both worlds.
@@ A.2 Derivation of Online Logistic Regression Laplace Approximation Update
This derivation was inspired by Ch. 4 of (Bishop and Nasrabadi, 2006)’s discussion on Laplace approximations and Bayesian logistic regression. The key idea is to Taylor expand the log-posterior around the current mean, which yields approximate closed-form update formulas while preserving approximate Gaussian conjugacy.
@@ A.3.2 Technical Assumptions/Conditions
Assumptions A.1-A.2 are the standard conditions under which a CLT for martingale difference arrays holds (c.f. Hall and Heyde (2014, Corollary 3.2)). Assumption A.3 is mild and is used to facilitate consistency of the variance estimator. Intuitively, in our setting of LLM evaluation with our proposed hybrid sampling policy (c.f. Eq. 10), the variance stabilization assumption is reasonable because as more outcome labels are observed, the model-factor estimate $(\hat{u}^{(t)},\hat{\Sigma}^{(t)})$ stabilizes, which stabilizes $\hat{p}^{(t)}$, and hence $q_{t}(\cdot)$, too (because $(\alpha_{t},\beta_{t})$ also stabilize). In addition, the Lindeberg condition is satisfied, by construction, under our hybrid sampling policy because $\tau$-mixing lower-bounds $q_{t}(j)$, avoiding extreme inverse-probability weights, and since $z_{j},\hat{p}^{(t)}_{j}\in[0,1]$, the increments $\Delta_{t}$ are uniformly bounded (by a constant depending only on $\tau$). The $\tau$-mixing also ensures the satisfaction of Assumption A.3.
@@ A.3.3 Asymptotic Normality via Martingale CLT
Under Assumptions A.1-A.2, the martingale central limit theorem (c.f.,  Hall and Heyde (2014, Corollary 3.2)) implies
@@
    Define $\tilde{\theta}^{(t-1)}:=\frac{1}{N_{q}}\sum_{j=1}^{N_{q}}\hat{p}^{(t-1)}_{j}$. Under Assumptions A.1A.2, Hall and Heyde (2014, Theorem 2.23) guarantees that
@@
    To show (17), we apply the Lenglart inequality (c.f., e.g., Whitt (2007, Lemma 3.7) or Jacod and Protter (2012, pp. 45)) to obtain